# Privacy Perspectives and Practices of Chinese Smart Home Product Teams

### v1

@@ 2.4 Legal/Regulatory Frameworks in China
P27-MS 25-34 Woman Master 1-5 yrs Guangdong Customer operations specialist 101-500 PE Smart thermostats, Smart lighting, Smart plugs, Kitchen appliances
@@
TABLE I: Participant demographics, and smart home device types that participants had designed and developed.
@@ 3 Methodology
We conducted semi-structured interviews with 27 Chinese smart home product team members from mainland China, using Mandarin for all interviews [^5] The translated screening survey, interview protocol, and codebook are available in this anonymous OSF repository. This section outlines our recruitment approach, participant demographics, interview protocol, data analysis details, ethical considerations, and study limitations.
@@ 3.1 Recruitment and Participants
We aimed to recruit a demographically-diverse sample of Chinese nationals, aged 18 or older, and employed as product team members at smart home companies based in mainland China. We identified participants with these job titles, using professional platforms (Maimai, LinkedIn), using our industry connections in Chinese smart home companies, and using snowball sampling [92]. We asked all interested individuals to complete a screening survey in Chinese. Based on industry standards and expected time commitment, we compensated each participant £50.
Our 27 participants, as detailed in Table I, included 11 women and 16 men; and 9 had a bachelor’s degree, 17 had a master’s degree, and 1 had a PhD. Participant roles included: 8 product/project managers (PM), 7 software/algorithm/hardware engineers (EN), 5 UX researchers/designers (UX), 4 legal privacy (data protection) officers/privacy engineers (LE), and 3 marketers/operations specialists (MS). Unlike previous studies that have focused on non-WEIRD technology (mainly software) who worked primarily in start-ups [65, 145, 95], most of our participants (20) were employed in large companies with more than 500 employees. Furthermore, 3 participants worked in medium-sized companies (101–500 employees), and 4 worked in small companies with fewer than 100 employees. We included participants from small to large companies to capture a broad spectrum of perspectives (see §5.2). Regarding company type, 8 participants were employed in state-owned enterprises (SOEs), 4 in multinational corporations (MNCs), and 15 in private enterprises (PEs).
Our participants were involved in designing and developing various smart home devices, including smart speakers (8), smart home cameras (16), door locks/doorbells (9), smart health devices/monitors (6), smart thermostats (8), baby cameras/monitors (5), smart lighting (6), smart TVs (4), smart plugs (7), kitchen appliances (7), smoke detectors/alarms (4), and smart home robots (3). Notably, only P2-PM and P27-MS expressed distrust toward the ownership of smart home devices, while 25 of them owned smart home devices in their own households.
@@ 3.2 Interview Procedure
Three of the authors (native Mandarin speakers) conducted all interviews remotely in Mandarin via VooV Meeting, with an average duration of 85.3 minutes. We audio-recorded and transcribed all sessions , and then we independently and manually checked the accuracy of all transcripts. The interviews began with participants outlining their roles within a product team, detailing the types of smart home devices they designed or developed, and their responsibilities. We then delved into their practices regarding user requirements gathering, focusing on UX and privacy-related requirements. We aimed to understand how they gathered the needs of different stakeholder groups, including at-risk groups (e.g., children, domestic workers) in multi-user smart homes, drawing inspiration from prior work like [98, 51, 20, 72, 24]. Additionally, we investigated how participants incorporated UX and privacy considerations into their design and development processes. We asked about data handling practices and legal compliance measures. To elicit deeper insights, we presented relevant examples from China (e.g., Didi case [^6] Cyberspace Administration of China (CAC) fined Didi \$1.2B for illegal collection and use of users’ personal data, violating laws on national data security and the protection of personal information.) to gauge their perspectives on technical interventions, such as data collection, transmission, deletion, and legal interventions like compliance training. These discussions aimed to uncover strategies for mitigating privacy concerns and risks in multi-user smart homes. We also sought their opinions on existing privacy laws and regulations in China (cf. §2.4).
@@ 3.3 Pilot Study
We carried out three pilot interviews to collect feedback. The pilots were not included in the final analysis. The feedback we obtained helped us revise some questions related to design and development aspects such as requirements gathering (e.g., asking about user involvement), collaboration (e.g., asking about collaborating with legal compliance and marketing team members), and adherence to legal regulations (e.g., probing into details about participants’ understanding of and compliance with PIPL).
@@ 3.4 Data Analysis
We achieved data saturation after 25 interviews and conducted an additional two interviews to verify this. Three authors checked after every few interviews whether data saturation had been reached, or whether to continue, noting insights began to repeat, and additional interviews did not provide new or significant variations to existing insights [43]. We inductively analyzed our interview transcripts. Initially, three authors familiarized themselves with the data, then each independently coded the same two randomly selected interview transcripts and created individual codebooks. They then met to discuss and resolve any disagreements, combining their individual codebooks into a merged one. After this, the three authors independently coded a different interview transcript to evaluate the merged codebook. They then met to discuss and resolve disagreements. After updating the codebook, they repeated this process three more times until no new codes emerged, indicating code saturation had been achieved. After they finalized the codebook, the first author coded all 27 interview transcripts, while the other two authors coded 13 and 14 transcripts, respectively, using the finalized codebook. We measured inter-rater reliability and checked agreement rates to ensure that coders used codes in a similar way, resulting in an average Cohen’s kappa of 0.88 for authors 1 and 2 and 0.92 for authors 1 and 3, indicating excellent agreement [39]. Finally, the three authors participated in organizing codes into themes specific to addressing our research questions. They also reviewed the excerpts on each topic to further refine the themes. We also recognize that our positionality, perspectives, and interpretations of the results are influenced by our diverse backgrounds and experiences (detailed in Appendix A).
@@ 3.5 Research Ethics
Our institution’s Research Ethics Committee reviewed and approved this study. Although we collected company names during screening to determine the size and type of organizations, we excluded them from our reported findings to protect the privacy of both participants and their companies. We used personal contact information (e.g., WeChat IDs) solely to communicate with potential participants during the recruitment process and securely deleted this data once recruitment was completed. After transcribing all interview recordings, we anonymized all transcripts by removing identifying details, such as names, company affiliations, and project-specific information. In addition, only researchers with access to the personal/contact details of the participants were involved in data collection and analysis.
@@ 3.6 Limitations
Given the qualitative nature of our study, the findings are not statistically generalizable beyond our purposive sample. However, our sample of participants spans a wide range of company types/sizes and roles within product teams in China’s smart home industry. This diversity enabled a rich and nuanced understanding of participants’ experiences, perspectives, and practices [108]. We also documented our study procedures in detail to support replication in other cultural and national contexts. All interviews were conducted and analyzed in Mandarin, with English translations provided by the three bilingual Chinese interviewers, helping to preserve the original meaning and minimizing potential loss of nuance. Although most of our participants (24) were aged 25-34 years, this aligns with broader demographic trends in China’s tech sector, where companies tend to favor younger employees for product roles, often perceiving those over 35 years of age as less efficient [78]. Lastly, as with all self-reported data, our findings may be subject to social desirability bias [88].
@@ 4 Results
In this section, we present the privacy perspectives and practices of participants, with a focus on multi-user smart home privacy in §4.1 (RQ1); the challenges faced by participants during the development life-cycle, as well as their views on power dynamics and privacy conflicts in smart homes in §4.2 (RQ2); and §4.3 describes the practices and strategies used by participants to address or mitigate privacy concerns of different stakeholder groups (RQ3). Because this research is qualitative, we did not attempt to quantify how many participants expressed a given view or theme; our research goal was to capture the richness and nuance of different experiences. We use terms like “most”, “many”, “some”, “several”, and “a few” only to give a rough insight into prevalence.
@@ 4.1 Perspectives, Experiences, and Practices
4.1.1 Privacy & security considerations during development in product teams
@@
Gathering security and privacy requirements. When we asked participants how they gathered user security and privacy-related requirements, most equated this process with complying with legal frameworks. These requirements were often tailored to the specific type of device being developed. For example, P5-UX highlighted the link between privacy and device type: “If the device is a surveillance camera or clearly collects user data, we will prioritize privacy and security.” Many participants regularly attended product team meetings centered on user privacy, particularly those in product management or legal/data compliance positions. For example, P12-PM underscored the necessity of complying with privacy laws, explaining: “I would definitely place security- and privacy-related requirements first. If a feature did not comply with privacy laws, I would ask my product team to redesign and launch it immediately.” P18-LE emphasized the importance of interpreting privacy-related regulations within the product team and described it as part of their responsibility: “Our team offers developers specialized legal and regulatory support by examining both domestic and international certification regulations and standards, providing comprehensive guidance to assist the development team with design and implementation.”
@@
Additionally, some participants had higher standards for safeguarding children’s privacy. For example, P3-PM highlighted that their product team thoughtfully balanced children’s privacy needs with parents’ caregiving requirements while designing and developing smart desk lamp devices equipped with camera sensors. P3-PM shared, “We gathered requirements from parents, and while most wanted constant monitoring features, we prioritized protecting children’s privacy over providing extensive monitoring capabilities to parents. Even video calls require children’s approval on their end.” Similarly, P12-PM mentioned that their product team implemented “higher privacy standards” to protect children’s data, deciding that data collected by smart baby monitors would only be streamed in real time and not stored. However, P17-EN noted the absence of specific data protection measures for children, explaining that from an algorithmic standpoint, there is no clear distinction between adult and child data with respect to privacy considerations.
@@
Prioritizing UX requirements gathering. Product teams, particularly product managers and UX/user researchers, employed various methods to collect requirements for the design and development of features and functionalities of smart home devices. They emphasized the importance of involving users to improve usability and functionality, as well as to identify practical issues during the initial requirements gathering and validation stages. In large product teams, some participants outsourced requirements gathering by asking third parties to conduct interviews and focus groups with users on their behalf; however, P5-UX noted that user research methods have limitations, as they often reflect individual viewpoints: “The biggest challenge is that the people we interview can only represent themselves—they cannot speak for all users.” Notably, marketing participants (P25–P27) had limited input on privacy and security, primarily due to minimal interaction with design and development teams. Even when probed, none reported involvement in communicating privacy features to users. Further, unlike typical software development, smart home device development often prioritizes hardware requirements (e.g., sensor integration). Some participants observed this shift, noting that it could result in privacy and security being de-prioritized. For example, P8-UX viewed privacy and security primarily as software issues rather than integral parts of the overall design process.
@@
Limited attention to privacy & security in testing. Many participants reported that they focused on assessing usability, stability, functionality, performance, and efficiency during the testing phase. Privacy and security were often neglected. In contrast, the teams of P9-UX, P10-PM, and P23-LE avoided involving external users to “protect trade secrets”, relying instead on internal testing conducted by product team members. Only P18-LE mentioned regular collaboration with the testing team on data compliance before product launch, stating: “We are the final checkpoint for the product. It must pass through us before reaching the market, making compliance essential. I see our role not as setting the upper limit but as defining the minimum standard.”
@@
Complying with applicable laws and regulations. Participants recognized the crucial role of legal compliance teams in shaping privacy practices. In addition to compliance experts, many participants, primarily product managers and senior leaders (e.g., software directors), noted that they frequently collaborated with legal compliance colleagues to assess security and privacy risks and ensure alignment with laws and regulations. As P12-PM noted: “Our legal team typically reviews the final designs for compliance before entering the development phase. They stays proactive by evaluating products whenever there are regulatory updates.” Some participants mentioned that their legal teams focused primarily on reviewing device privacy policies but did not always verify whether privacy requirements were fully met. P14-EN also expressed concerns about complying with new laws or regulations: “I know some small companies might rush to launch smart home apps before new laws take effect, trying to dodge compliance by exploiting regulatory loopholes. They focus more on getting to market first than on adhering to the upcoming regulations.”
@@
In addition, we observed that participants with over 10 years of experience and in managerial roles (e.g., P5-UX, P15-EN, & P24-EN) demonstrated greater awareness of privacy and data protection laws compared to less experienced participants. They advised their product teams to adjust feature designs to comply with legal requirements. For example, P5-UX altered the user consent dialog interface on their mobile app for smart home cameras to align with POPCPI regulations [^7] POPCPI’s articles 8 & 9 require data processors to protect children’s personal data and obtain consent from children’s guardians to process it.: “The law requires us to obtain consent from parents or guardians, so we implemented a dedicated consent option specifically for them. We also included a ‘refusal’ option in case the guardian does not agree to provide authorization.”
@@
Summary. Most participants equated gathering user security and privacy requirements with complying with Chinese privacy laws that prioritized national security over individual privacy; see §2.4. In contrast, gathering UX requirements—covering functionality and performance—involved active user engagement, often outsourced to third parties. Privacy and security received limited focus during testing.

### v2

@@ 2.4 Legal and Regulatory Frameworks in China
P27-MS 25-34 Woman Master’s 1-5 yrs Guangdong Customer operations specialist 101-500 PE Smart thermostats, Smart lighting, Smart plugs, Kitchen appliances
@@
TABLE I: Participant demographics and the types of smart home devices they had designed/developed.
@@ 3 Methodology
We conducted semi-structured interviews with 27 Chinese smart home product team members from mainland China, using Mandarin for all interviews [^5] The translated screening survey, interview script, and codebook, as well as a comparative table of PIPL, DSL, CSL, and POPCPI, are available in this OSF repository.. This section outlines our recruitment approach, participant demographics, interview protocol, data analysis details, ethical considerations, and study limitations.
@@ 3.1 Recruitment and Participants
We aimed to recruit a diverse sample of Chinese nationals who were 18 years or older and were employed as product team members in smart home companies based in mainland China, ensuring variation in roles, company sizes, and experiences. We identified participants with these job titles using professional platforms (Maimai, LinkedIn), our industry connections with Chinese smart home companies, and snowball sampling, in which participants referred peers in different roles, either within their own company or from other organizations [80]. We asked all interested individuals to complete a screening survey in Chinese. Based on industry standards and expected time commitment, we compensated each participant £50.
Our 27 participants, as detailed in Table I, included 11 women and 16 men; and 9 had a bachelor’s degree, 17 had a master’s degree [^6] In China’s IT sector, it is common for product team members to hold a master’s degree [82]., and 1 had a PhD. Participant roles included: 8 product/project managers (PM), 7 software/algorithm/hardware engineers (EN), 5 UX researchers/designers (UX), 4 legal privacy (data protection) officers/privacy engineers (LE), and 3 marketers/operations specialists (MS). Although we recruited some participants with company-specific job titles, these titles could be broadly categorized. For example, “algorithm engineer” falls under the broader category of “software engineer.” The responses from participants with such titles were consistent with the overall insights provided by other participants in the same broader category and did not influence the final themes. Unlike previous studies that have focused on non-WEIRD technologies (mainly software systems) and participants who worked primarily in start-ups [58, 130, 83], most of our participants (20) worked in large companies with more than 500 employees. Furthermore, 3 participants worked in medium-sized companies (101–500 employees), and 4 worked in small companies with fewer than 100 employees. We included participants from small to large companies to capture a broad spectrum of perspectives (§5.2). Regarding company type, 8 participants were employed in state-owned enterprises (SOEs), 4 in multinational corporations (MNCs), and 15 in private enterprises (PEs). 22 participants were employed by different companies, while 5 worked in two firms: 2 in one (P9, P24) and 3 in another (P11, P13, P23). In addition, 4 employed by MNCs worked on devices for the Chinese market and operated under local compliance requirements.
Our participants were involved in designing and developing various smart home devices, including smart speakers (8), smart home cameras (16), door locks/doorbells (9), smart health devices/monitors (6), smart thermostats (8), baby cameras/monitors (5), smart light bulbs (6), smart TVs (4), smart plugs (7), kitchen appliances (7), smoke detectors/alarms (4), and smart home robots (3). Notably, only P2-PM and P27-MS expressed distrust regarding the ownership of smart home devices, while the remaining 25 participants owned such devices in their own households.
@@ 3.2 Interview Procedure
Authors 1-3 are native Mandarin speakers fluent in English. Author 1 conducted all interviews remotely in Mandarin via VooV Meeting, while Authors 2 and 3 attended to take notes. The interviews had an average duration of 85.3 mins, ranging from 44.5 to 112 mins. We audio-recorded and transcribed all sessions, and then we independently and manually checked the accuracy of all transcripts. The interviews began with participants outlining their roles within a product team, detailing the types of smart home devices they designed or developed, and their responsibilities. We then asked questions about participants’ privacy and/or legal obligations. Although these questions were introduced early in the guide—after asking about roles, team structure, and responsibilities to build rapport—they were intended to examine participants’ roles in data governance. The questions were broadly framed to avoid assumptions about participants’ views and were followed by exploratory probes. We then investigated participants’ practices regarding the collection of user requirements, with a focus on UX and privacy-related requirements. We aimed to understand how participants examined the needs and privacy concerns of different stakeholder groups, including at-risk groups (e.g., children, domestic workers) in multi-user smart homes, drawing on previous work such as [85, 47, 18, 64, 25]. Further, we examined how participants incorporated UX and privacy considerations throughout their design and development processes. It is worth noting that questions on UX requirements gathering (including privacy concerns) were directed primarily at participants in roles responsible for this task (e.g., product managers), while other participants (e.g., compliance officers) were asked about deriving requirements from laws, regulations, or sector-specific guidelines.
To reduce social desirability bias, we began with neutral, open-ended questions (e.g., “What are the main challenges that you and your product team face when designing and/or developing smart home devices?”) and emphasized that there were no right or wrong answers. We also asked about data handling practices and legal compliance measures. To elicit deeper insights, we presented relevant examples from China (e.g., Didi case [^7] The Cyberspace Administration of China (CAC) fined Didi \$1.2 billion for the illegal collection and use of users’ personal data, in violation of laws on national data security and personal information protection.) to gauge their perspectives on technical interventions such as data collection, transmission, and deletion, as well as legal interventions like compliance training. These real-world examples encouraged participants to reflect on and propose technical, social, or legal interventions. We also sought their opinions on existing privacy laws and regulations in China (cf. §2.4).
@@ 3.3 Pilot Study
We carried out three pilots to collect feedback. The pilots are not included in the final analysis. The feedback we obtained helped us revise some questions related to design and development aspects such as requirements gathering (e.g., asking about user involvement), collaboration (e.g., asking about collaborating with legal compliance and marketing team members), and adherence to legal regulations (e.g., probing into details about participants’ understanding of and compliance with PIPL). We also reflected critically on how the phrasing, sequencing, and specificity of questions might influence participants’ responses. For example, we reworded compound or overly narrow questions to ensure participants in various roles could respond meaningfully.
@@ 3.4 Data Analysis
Sample size in qualitative research remains a topic of debate [20], particularly given the inherently subjective nature of determining data saturation—commonly defined as the point at which no new data is found [21]. In our study, we achieved data saturation after 25 interviews and conducted two additional interviews to confirm this. Specifically, three authors reviewed the data after every few interviews to assess whether saturation had been reached or if further interviews were needed, noting that insights began to repeat and that additional interviews did not provide new or significant variations to the existing data [41]. Our sample size also aligns with widely accepted qualitative research standards, which prioritize rich, contextualized understanding over statistical generalizability [108].
We inductively analyzed our interview transcripts using in-vivo coding to remain close to participants’ perspectives. Authors 1–3 coded and analyzed all transcripts in Chinese to preserve meaning and prevent potential loss that could have occurred if the transcripts had been translated into English first. Specifically, Authors 1–3 independently coded three randomly selected transcripts to develop an initial codebook. The three authors then met and, through joint discussions, identified codes, refined overlapping codes, clarified boundaries (e.g., distinguishing “Prioritize privacy & security requirements” from “Data management mechanisms”), and established rules for the consistent application of codes throughout the coding process. After resolving disagreements, they merged their individual codebooks into a single codebook. Using the merged codebook, they randomly selected and independently coded an additional transcript (different from the previous three), updating the codebook as needed. This process was repeated two more times to finalize the codebook and achieve code saturation.
After finalizing the codebook, each transcript was independently coded by two researchers (Interviews 1–13 by Authors 1 and 2, and 14–27 by Authors 1 and 3) to enhance rigor and reliability, prevent individual researcher bias, capture nuanced interpretations, and ensure consistent application of the codebook [74]. We measured inter-rater reliability and checked agreement rates, resulting in an average Cohen’s kappa of 0.88 for Authors 1 and 2 and 0.92 for Authors 1 and 3, indicating excellent agreement [37].
As a research team, we then iteratively identified, developed, and refined themes until they addressed our research questions, aligning with the 20 themes presented in §4. In particular, we organized codes into themes specific to our research questions and reviewed the corresponding excerpts for each topic to further refine the themes. Finally, Author 1 translated the codebook and participant quotes reported in §4 into English. These translations were subsequently reviewed and verified by Authors 2 and 3 to ensure accuracy and minimize potential translation gaps [119].
@@ 3.5 Research Ethics
The Research Ethics Committee of King’s College London reviewed and approved this study (ID: LRS/DP-22/23-35814). While company names were collected during screening to assess the size and type of organizations, they were excluded from our reported findings to protect the privacy of both participants and their companies. Personal contact information (e.g., WeChat IDs) was used solely to communicate with potential participants during recruitment and was securely deleted once recruitment was completed. Following the transcription of all interview recordings, all transcripts were anonymized by removing identifying details, including names, company affiliations, and project-specific information. To safeguard participant confidentiality, only Author 1 retained the contact details of participants.
3.6 Author Positionality
Since privacy can vary significantly based on cultural, social, and individual factors, we did not provide participants with a specific definition of privacy, as existing definitions in the academic literature are predominantly shaped by WEIRD values. Instead, we explored participants’ privacy perspectives and practices as influenced by their backgrounds and experiences, given that our study focuses on a non-WEIRD country. Although participants did not explicitly mention the term “Confucianism” during the interviews, they discussed values aligned with it, such as filial piety (§4.2.2). Recognizing these culturally ingrained themes, we incorporated Confucianism into our analysis and reporting of results. Rather than treating it as a discrete code in our codebook, Confucianism emerged as an overarching theme during the process of developing and refining our themes.
In addition, our positionality and interpretation of the results were shaped by our backgrounds and experiences. All authors are expert researchers in human-centered computing, with extensive experience working with both majority populations and at-risk groups. Although all authors are trained researchers based primarily in Western institutions, the interviewers and data analysts contributed additional perspectives rooted in their personal backgrounds, having been born and raised in China. This dual perspective allowed them to approach the interviews and interpret the data with both academic rigor and deep cultural understanding. Also, Author 4 specializes in technology privacy law and public policy in both Chinese and Western contexts, including the UK and the US.
@@ 3.7 Limitations
Given the qualitative nature of our study, the findings are not statistically generalizable beyond our sample. However, our sample of participants spans a wide range of company types/sizes and roles within product teams in China’s smart home industry. This diversity enabled a rich and nuanced understanding of participants’ experiences, perspectives, and practices [94]. We also documented our study procedures in detail to support replication in other cultural and national contexts. Although most of our participants (24) were between 25 and 34 years old, this reflects broader demographic trends in China’s tech sector, where companies often favor younger employees for product roles and view those over 35 years of age as less efficient [69]. Furthermore, we recognize that the interview guide might have included some leading questions, which could have affected participants’ responses, despite our efforts to reduce bias by using open-ended and role-specific phrasing (cf. §3.2). Lastly, as with all self-reported data, our findings might be influenced by social desirability bias [77].
@@ 4 Results
In this section, we present the privacy perspectives and practices of participants, with a focus on multi-user smart home privacy in §4.1 (RQ1); the challenges faced by participants during the development life-cycle, as well as their views on power dynamics and privacy conflicts in smart homes in §4.2 (RQ2); and §4.3 describes the practices and strategies used by participants to address or mitigate privacy concerns of different stakeholder groups (RQ3). Because this research is qualitative, we did not attempt to quantify how many participants expressed a given view or theme; our research goal was to capture the richness and nuance of different experiences. We use terms like “most”, “many”, “some”, “several”, and “a few” only to give a rough insight into prevalence. Across the sections, we treat privacy as contextual and selectively implemented, with protections shaped by legal mandates (e.g., POPCPI) and by cultural or dignity-based values (e.g., respectful blurring in elder care).
@@ 4.1 Perspectives, Experiences, and Practices
4.1.1 Privacy & security considerations during development (product teams)
@@
Gathering security and privacy requirements. When we asked participants how they gathered user security and privacy-related requirements, most equated this process with complying with legal frameworks. These requirements were often tailored to the specific type of device being developed. For example, P5-UX highlighted the link between privacy and device type: “If the device is a surveillance camera or clearly collects user data, we will prioritize privacy and security.” Many participants regularly attended product team meetings centered on user privacy, particularly those in product management or legal compliance positions. For example, P12-PM underscored the necessity of complying with privacy laws, explaining: “I would definitely place security- and privacy-related requirements first. If a feature did not comply with privacy laws, I would ask my product team to redesign and launch it immediately.” P18-LE emphasized the importance of interpreting privacy-related regulations within the product team and described it as part of their responsibility: “Our team offers developers specialized legal and regulatory support by examining both domestic and international certification regulations and standards, providing comprehensive guidance to assist the development team with design and implementation.”
@@
Additionally, some participants described applying protections selectively and contextually, guided by both legal mandates and considerations of dignity. For example, P3-PM highlighted that their product team thoughtfully balanced children’s privacy needs with parents’ caregiving requirements while designing and developing smart desk lamp devices equipped with camera sensors. P3-PM shared, “We gathered requirements from parents, and while most wanted constant monitoring features, we prioritized protecting children’s privacy over providing extensive monitoring capabilities to parents. Even video calls require children’s approval on their end.” Participants explicitly connected these choices to legal obligations under POPCPI for children’s data and to culturally-grounded commitments to dignity-respecting design. Several noted that they maintained these protections even when they believed that parents favored continuous monitoring. Similarly, P12-PM mentioned that their product team implemented “higher privacy standards” to protect children’s data, deciding that data collected by smart baby monitors would only be streamed in real time and not stored. These decisions were presented as a negotiation between perceived user attitudes, relevant legal frameworks (e.g., POPCPI), and broader cultural expectations. However, P17-EN noted the absence of specific data protection measures for children, explaining that from an algorithmic standpoint, there is no clear distinction between adult and child data with respect to privacy considerations.
@@
Prioritizing UX requirements gathering. Product teams, particularly product managers and UX/user researchers, employed various methods to collect requirements for the design and development of features and functionalities of smart home devices. They emphasized the importance of involving users to improve usability and functionality, as well as to identify practical issues during the initial requirements gathering and validation stages. In large product teams, some participants outsourced requirements gathering by asking third parties to conduct interviews and focus groups with users on their behalf; however, P5-UX criticized the industry’s reliance on convenience sampling, arguing that it limited research findings to individual perspectives shaped by narrow or biased participant pools: “The biggest challenge is that the people we interview can only represent themselves—they cannot speak for all users.” Notably, privacy communication was managed by legal officers or product managers rather than UX/user researchers or marketing staff. Marketing participants (P25–P27) reported having limited input on privacy and security, largely because of minimal interaction with design and development teams. Even when probed, none reported involvement in communicating privacy features to users. Further, unlike typical software development, smart home device development often prioritizes hardware requirements (e.g., sensor integration). Some participants observed this shift, noting that it could result in privacy and security being de-prioritized. For example, P8-UX viewed privacy and security primarily as software issues rather than integral parts of the overall design process.
@@
Limited attention to privacy & security in testing. Many participants reported that they focused on assessing usability, stability, functionality, performance, and efficiency during the testing phase. Privacy and security were often neglected. In contrast, the teams of P9-UX, P10-PM, and P23-LE avoided involving external users to “protect trade secrets”—relying instead on internal testing conducted by product team members. Only P18-LE mentioned regular collaboration with the testing team on data compliance before product launch, stating: “We are the final checkpoint for the product. It must pass through us before reaching the market, making compliance essential. I see our role not as setting the upper limit but as defining the minimum standard.”
@@
Complying with applicable laws and regulations. Participants recognized the crucial role of legal compliance teams in shaping privacy practices. In addition to compliance experts, many participants, primarily product managers and senior leaders (e.g., software directors), noted that they frequently collaborated with legal compliance colleagues to assess security and privacy risks and ensure alignment with laws and regulations. As P12-PM noted: “Our legal team typically reviews the final designs for compliance before entering the development phase. They stay proactive by evaluating products whenever there are regulatory updates.” Some participants mentioned that their legal teams focused primarily on reviewing device privacy policies but did not always verify whether privacy requirements were fully met. P14-EN also expressed concerns about complying with new laws or regulations: “I know some small companies might rush to launch smart home apps before new laws take effect, trying to dodge compliance by exploiting regulatory loopholes. They focus more on getting to market first than on adhering to the upcoming regulations.”
@@
In addition, we observed that participants with over ten years of experience and in managerial roles (e.g., P5-UX, P15-EN, & P24-EN) demonstrated greater awareness of privacy and data protection laws compared to less experienced participants. They advised their product teams to adjust feature designs to comply with legal requirements. For example, P5-UX altered the user consent dialog interface on their mobile app for smart home cameras to align with POPCPI regulations [^8] Under POPCPI Articles 8 and 9, data processors are required to safeguard children’s personal data and obtain consent from their legal guardians prior to processing.: “The law requires us to obtain consent from parents or guardians, so we implemented a dedicated consent option specifically for them. We also included a ‘refusal’ option in case the guardian does not agree to provide authorization.”
@@
Summary. Most participants equated gathering privacy & security requirements with complying with Chinese privacy laws that prioritized national security over individual privacy; see §2.4. In contrast, gathering UX requirements—covering functionality and performance—involved active user engagement, often outsourced to third parties. Privacy was applied selectively and contextually—driven by (i) legal requirements (e.g., POPCPI for children’s data) and (ii) cultural or dignity values (e.g., respectful blurring in elder-care scenarios). Privacy and security received limited focus during testing.